govt. data shows that 10% of males under age 25 are unemployed. A random sample is taken of 400 males who are in the labor force and under age 25. Find the chance that the sample unemployment rate is 12% or more.

Answer by stanbon(75887)

govt. data shows that 10% of males under age 25 are unemployed. A random sample is taken of 400 males who are in the labor force and under age 25. Find the chance that the sample unemployment rate is 12% or more.
-----
z(0.12) = (0.12-0.10)/sqrt[0.10*0.90/100] = 0.667
----
P(p >= 0.12) = P(p >= 0.667) = 0.2525
=====================================
Cheers,
Stan H.